# Language Models Learn Rare Phenomena from Less Rare Phenomena: The Case of the Missing AANNs

### v1

@@ 6 Conclusion
There is increasing interest in computational linguistics in how language models can handle what has been variously called the “long tail” of language (Prange et al., 2021), “extremely rare constructions” (Potts, 2023), “exceptions to syntactic rules” (Leong and Linzen, 2023), “rare linguistic phenomena” (Weissweiler et al., 2024), inter alia. Studies of such phenomena are important first because LLMs (and LMs and statistical models in general) are known to be extremely sensitive to frequency and to perform far better in data-rich environments and, second, because the human ability to generalize to rare phenomena is central to knowledge of language.
We found that LMs trained on a human-scale of data can learn a rare construction like the aann. We found that this learning occurs even without veridical instances of the aann construction in the training data, and that it is mediated by occurrences of other related constructions in training. As such, these results join a growing body of work showing the ability of LLMs to learn constructions (Tayyar Madabushi et al., 2020; Tseng et al., 2022; Li et al., 2022; Veenboer and Bloem, 2023).
@@ 7 Limitations
In future work, it would be valuable to extend this method to a wider range of constructions. But scaling this approach up is not straightforward since it requires identifying and extracting idiosyncratic constructions, and—more onerously—developing testable hypotheses about what makes them learnable from limited amounts of data. While this is a limitation, it also calls for more synergistic collaborations between theoretical and computational linguists.
Another limitation is that our method requires repeated training of LMs from scratch which can be computationally expensive. Alternate methods could be to ablate knowledge of particular hypotheses using representational editing methods such as AlterRep (Ravfogel et al., 2021), etc.
Unlike Weissweiler et al. (2022), we do not test the ability to interpret these constructions for downstream tasks. Instead, we focus on linguistic form. Extending these results to semantic tasks would be informative.
@@ 8 Acknowledgments
(KM)2 acknowledge funding from NSF Grant 2104995 awarded to Kyle Mahowald. For helpful conversations, we thank Adele Goldberg, Leonie Weissweiler, the computational linguistics research group at UT Austin, the syntax-semantics research group at UT Austin, the audience at the Texas Linguistics Society meeting. We thank Chris Potts for his paper on the PiPP construction (Potts, 2023) which inspired the “keys to all of this” idea.
Appendix A LM training details
As mentioned in the main text (see Section 2), we use the OPT architecture (Zhang et al., 2022) to train our LMs on all versions of the BabyLM corpus. This was the best performing autoregressive LM in the BabyLM Competition (Warstadt et al., 2023). For each instance of the BabyLM (ablated or otherwise), we tune the learning rate [^5] We searched the following set: {1e-4, 3e-4, 1e-3, 3e-3} based on the validation set, and use the best learning rate as a result of the tuning to train an additional two language models using different seeds. As a result, for each ablation of the BabyLM corpus, we run 6 LM training experiments, which amounts to a whopping 90 LMs for all our experiments. Table 3 contains further details of the training.
@@ Appendix A LM training details
Appendix B Detecting aanns and related phenomena
In this section, we briefly describe our methods to extract constructions and phenomena relevant to this paper from the BabyLM corpus (Warstadt et al., 2023). Our methods primarily rely on: 1) the surface form of the sentences in the corpus; 2) their corresponding part-of-speech (POS) tag sequences; and in a few cases, 3) their depedency parses. For the latter two, we used spacy (Honnibal et al., 2020), specifically, its en_web_trf model, which is based on the RoBERTa-base LM (Liu et al., 2019). Next we describe how we used these artifacts to detect our target constructions:
B.1 AANNs
To detect AANNs we primarily rely on POS-tagged sequences, and construct a regex pattern over them [^6] In reality this was constructed over several iterations, taking into account many different possible realizations of the construction in free text. which is able to robustly detect aanns:
@@ B.1 AANNs
where we restrict the determiner (DT) to be either ‘a’, ‘an’, or ‘another’. This regex permits multiple adjectives (an exhilarating and marvelous three months) optional adverbs (an excruciatingly painful two semesters), multi-word noun phrases with plural head-nouns (a refreshing two glasses of aperol spritz), numeral-expressions involving subordinate clauses (a measly three to five days), among other potential edge cases. We additionally use the following adjectives as proxies for numerals, as per the guidelines of Kayne (2007) and Solt (2007):
numeral_proxies = [’few’, ’dozen’, ’couple’, ’several’, ’many’, ’more’]
For instance, the following examples would all count as instances of the aann:
@@ B.2 DT ANNs
B.3 A few/couple/dozen NOUNs
@@
B.4 Measure NNS with Singular Verbs
@@
Appendix C A/An + ADJ/NUM frequency balancing

### v2

@@ 6 Conclusion
Theoretically, there is, for good reason, considerable interest in how language models can handle what has been variously called the “long tail” of language (Prange et al., 2021), “extremely rare constructions” (Potts, 2023), “exceptions to syntactic rules” (Leong and Linzen, 2023), “rare linguistic phenomena” (Weissweiler et al., 2024), inter alia. Studies of such phenomena are important first because LMs (and statistical models in general) are sensitive to frequency and often perform far better in data-rich environments and, second, because the human ability to generalize to rare phenomena is central to linguistics.
Empirically, we found that LMs trained on modest amounts of data can learn a rare construction like the aann. We found that this learning occurs even without veridical instances of the construction in the training data, and that it is mediated by occurrences of other related constructions in training. As such, these results join a body of work showing the ability of LLMs to learn rare phenomena (Tayyar Madabushi et al., 2020; Tseng et al., 2022; Li et al., 2022; Veenboer and Bloem, 2023) and to generalize from limited data in meaningful ways.
Methodologically, this work leave us optimistic that “controlled rearing” of LMs is a fecund method for understanding models, as well as for gleaning insight into human language more generally.
@@ 7 Limitations
In future work, it would be valuable to extend this method to a wider range of constructions. But scaling this approach up is not straightforward since it requires identifying and extracting idiosyncratic constructions, and—more onerously—developing testable hypotheses about what makes them learnable from limited amounts of data. Future work will likely benefit from synergistic collaborations between theoretical and computational linguists.
Another limitation is that our method requires repeated training of LMs from scratch which can be computationally expensive. Alternate methods could be to ablate knowledge of particular hypotheses using representational editing methods, though these may not guarantee perfect removal of the knowledge of targeted constructions.
Unlike Weissweiler et al. (2022), we do not test the ability to interpret these constructions for downstream tasks. Instead, our ablations target linguistic form alone, and preliminary experiments suggest that our ablations and manipulations leave the lexical semantic properties of the aann unchanged (see App. E). Extending our ablation method to target these properties more directly would be quite informative.
Finally, this work only studies a rare construction in English, and on LMs that are trained on English text data. While this is a limitation of the paper, the paradigm introduced can be readily used in future work to study hypotheses and perform indirect evidence elicitation in multi-lingual LMs.
@@ 8 Acknowledgments
(KM)2 acknowledge funding from NSF Grant 2104995 awarded to Kyle Mahowald. For helpful conversations, we thank Adele Goldberg, Leonie Weissweiler, Nathan Schneider, Tom McCoy, the computational linguistics research group at UT Austin, the syntax-semantics research group at UT Austin, audiences at the Texas Linguistics Society meeting, Edinburgh University Department of Linguistics, University of Antwerp CLiPS group, attendees of the ANN workshop in Amsterdam, and the Brown University language group. We thank Lisa Bylinina for exceptionally helpful comments on an earlier draft. We thank Chris Potts for his paper on the PiPP construction (Potts, 2023) which inspired the “keys to all of this” idea.
Appendix A Dataset Access and Licensing
The aann acceptability dataset by Mahowald (2023) is released using the MIT License and was accessed from the author’s public github repo. [^1] https://github.com/mahowak/aann-public The BabyLM dataset222accessed from https://babylm.github.io/ does not have a single license of its own but instead inherits the licenses of its constituents: CHILDES (MacWhinney, 2000), BNC Dialogue portion, [^3] http://www.natcorp.ox.ac.uk Children’s Book Test (Hill et al., 2016), Children’s Stories Text Corpus, [^4] https://www.kaggle.com/datasets/edenbd/children-stories-text-corpus Standardized Project Gutenberg Corpus (Gerlach and Font-Clos, 2020), OpenSubtitles (Lison and Tiedemann, 2016), QCRI Educational Domain Corpus (Abdelali et al., 2014), Wikipedia, [^5] https://dumps.wikimedia.org/enwiki/20221220/ Simple Wikipedia, [^6] https://dumps.wikimedia.org/simplewiki/20221201/ Switchboard Dialog Act Corpus (Stolcke et al., 2000). Since this dataset was specifically released to train LMs, we work under the assumption that our LMs do not violate its license policies. We will follow the inherited licenses’ policies while making the trained LMs and ablated BabyLM data public, and refrain from releasing them if we find them to be in violation of the policies.
Appendix B LM training details
As mentioned in the main text (see § 2), we use the OPT architecture (Zhang et al., 2022) to train our LMs on all versions of the BabyLM corpus. This was the best performing autoregressive LM in the BabyLM Competition (Warstadt et al., 2023). For each instance of the BabyLM (ablated or otherwise), we tune the learning rate [^7] We searched the following set: {1e-4, 3e-4, 1e-3, 3e-3} based on the validation set, and use the best learning rate as a result of the tuning to train an additional two language models using different seeds. As a result, for each ablation of the BabyLM corpus, we run 6 LM training experiments, which amounts to a whopping 114 LMs for all our experiments. Table 3 contains further details of the training.
@@ Appendix B LM training details
Appendix C Detecting aanns and related phenomena
In this section, we briefly describe our methods to extract constructions and phenomena relevant to this paper from the BabyLM corpus (Warstadt et al., 2023). Our methods primarily rely on: 1) the surface form of the sentences in the corpus; 2) their corresponding part-of-speech (POS) tag sequences; and in a few cases, 3) their dependency parses. For the latter two, we used spacy (Honnibal et al., 2020), specifically, its en_web_trf model, which is based on the RoBERTa-base LM (Liu et al., 2019). Next we describe how we used these artifacts to detect our target constructions:
C.1 aanns
To detect aanns, we constructed a regex-based pattern-matcher which operated over a POS-tagged version of the BabyLM corpus. We started with an initial regex pattern (Regex v1), as shown in LABEL:regexv1:
Figure 6: Pipeline to assess the recall of our aann-detecting regex patterns, along with examples of cases missed by each regex. The recall for our final regex (Regex v3) is 95% (missing only one instance where there was a typo), and it is able to handle complex and sophisticated forms of the construction.
Listing 1: Regex v1.
@@ Appendix C Detecting aanns and related phenomena
here we restrict the determiner (DT) to be either ‘a’, ‘an’, or ‘another’. This regex permits multiple adjectives (an exhilarating and marvelous three months) optional adverbs (an excruciatingly painful two semesters), multi-word noun phrases with plural head-nouns (a refreshing two glasses of aperol spritz), numeral-expressions involving subordinate clauses (a measly three to five days), among other potential edge cases.
We then tested this regex pattern on a large sample of utterances which we extracted using a permissive regex applied to the 10M-token version of BabyLM (a subset of our 100M training set), which looked for any “a” or “an” or “another” that appeared sequentially prior to a numeral as well as a plural noun in a sentence. Importantly this regex filter did not rely on any POS tagging, to avoid issues attributable to tagging errors. We hand-annotated a sample of 3000 utterances from this set, and found 49 legitimate aanns. [^8] In reality, we found 50, but rejected one of them: “a good 1-2” of snow…, where ‘”’ is inches. This would have never been caught unless we are to include ‘”’ in our pipeline which would conflate other uses of quotes. Our Regex v1 only detected 29 of these, meaning its recall was around 59%.
We then developed a second version of the regex (Regex v2; see LABEL:regexv2) to handle cases that the above regex pattern missed (e.g., using participle modifiers, occurrence of punctuation or extra spaces in between, accounting for hedging, a case where ‘record’ was used as a modifier, etc.).
Listing 2: Regex v2.
[…]
To test Regex v2, we again used the permissive regex and extracted an additional 1000 samples from our training set. On hand-annotating them, we found 24 valid aanns, out of which Regex v2 detected 18, bringing up the recall to 75%.
In both the previous cases, we were post-processing the detected aanns to include certain adjectives (few, dozen, couple, several, many, more) as numerals, as per the guidelines of Kayne (2007) and Solt (2007). This allows the following to also be considered instances of the aann:
@@
At the same time, this also ends up including cases such as:
[…]
Similarly, we had to include NN within our adjective span of the regex pattern to accommodate ‘record’ when used as/as part of a modifier (e.g., a record-high 60 miles per hour), but this exploded the number of “detected” aanns, lowering our precision drastically, due to which we omitted it.
To address these issues, we decided to pre-process the POS-tagged corpora prior to using our regex, where we substituted articles of interest with the ‘ARTICLE’ tag, substituted record when preceeded by an article with the ‘RECORD’ tag, and numeral proxies with the ‘FEW’ tag, though ensuring that it appeared linearly after a known adjective which was not a numeral proxy. This led to the creation of Regex v3 (LABEL:finalregex):
Listing 3: Regex v3 (final). Tags such as ARTICLE, RECORD, FEW are added after POS-tagging to include certain special tokens.
[…]
This was able to handle the idiosyncracies of all previously detected aanns. We again extracted a further additional 1000 samples to hand-annotate and found 18 attested aanns. Regex v3 was able to detect 17 out of these (recall of 95%), missing out on only one where an incorrect form was used in lieu of a plural noun (e.g., pound instead of pounds). We don’t really consider this a meaningful missed example since the singular noun actually makes this a degenerate aann, not a genuine one (but, to be conservative, count it as a miss for assessing a worst-case recall estimate). At this point, we stopped further refining our regex and used Regex V3 as our final detector, while also acknowledging that it is perhaps impossible to guarantee whether every single aann instance is captured by the regex. Fig. 6 shows our recall analysis pipeline in a nutshell.
@@ C.2 DT ANNs
C.3 A few/couple/dozen NOUNs
@@
C.4 Measure NNS with Singular Verbs
@@
Appendix D A/An + ADJ/NUM frequency balancing
@@
Appendix E Lexical semantic constraints on aann slots
Fig. 7 shows the breakdown of acceptability ratings from humans and LMs across various adjective and noun classes.
Figure 7: z-scored aann acceptability ratings elicited from Humans (scale of 1-10) and LMs (slors) trained on corpora with (1) aanns removed (i.e., no aann); and (2) left unablated. Ratings broken down based on adjective and noun classes. Ratings are computed for each system based on Mahowald (2023)’s entire dataset, which consists of human derived acceptability judgments on 3,420 different types of aanns.
Appendix F Variability Analysis
In § 5 we compared aann-generalization of LMs trained on BabyLM versions which differed in the amount of variability that was present in the aanns that the models were exposed to. In particular, we operationalized variability in terms of the slot-fillers of the adjective/numeral/noun slots, both together as well as individually. Table 4 shows three examples of high and low variability items (each) for the four different slot-filler based considerations in our experiments.
Slot High Variability Low Variability
Instance Freq. Instance Freq.
All impressive 30 appearances 1 great many things 42
massive 108 years 1 good many years 21
reported 14 million dolars 1 additional two years 4
Adj career-high 38 great 355
staggering 12 additional 111
measly 7 mere 60
Num 20 32 two 174
couple 17 five 67
seven to eight 1 few 64
Noun dollars 15 years 254
students 8 miles 77
kangaroos 1 hours 42
Table 4: Examples of slot fillers that were ablated as part of our variability experiments, along with their frequency in the training data, across all slots considered (All open slots, Adjective-only, Numeral-only, and Noun-only).